OpenStudy (anonymous):

7/9 * 1/2 - 1/6 simplest form

OpenStudy (anonymous):

@lizz123

OpenStudy (lizz123):

Multiply first. \[\frac{ 7 }{ 9 } \times \frac{ 1 }{ 2 } = \frac{ 7 }{ 18 }\]

OpenStudy (lizz123):

But you still need to subtract

OpenStudy (anonymous):

6/12

OpenStudy (lizz123):

which it can be simplified

OpenStudy (lizz123):

by 3

OpenStudy (anonymous):

1/2

OpenStudy (lizz123):

yeah that is what I got too

OpenStudy (anonymous):

its not it

OpenStudy (lizz123):

wait it is 2/9